Make New Entry,
Make Followup Entry
User name dalton
Log entry time 11:34:25 on November 27, 2009
Entry number 303020
This entry is a followup to: 302755
keyword=Run coordinator daily summary
Run Coordinator's Report for Thursday November 26 Day to Friday
November 27 Owl
- 4 Owl shifts in the next week need a shift leader.
Please sign up!
- A controlled access was taken during during the Day shift to adjust
some cable delays which resulted in a significant improvement in
- We took 16.5 hours of ABU over the last 24 hours in which we
completed slug 8 and achieved ~400 K good LD2 pairs for slug 9.
- High current production running throughout weekend.
- 4 Owl shifts in the next week need a shift leader. Please
- Shift Leaders: please write a detailed shift summary (including
runs taken and any
activities) and submit the BTA for each shift. Please log CODA reboots
in the halog and shift summary.
- 3rd person: Please make sure to start and stop the HRS DAQ
at the same time as the parity DAQ and to do the prompt analysis and the
HRS analyzer for every run.
- 3rd person: please keep a careful eye on the online data display.
Most problems can be caught quickly this way.
- Please make sure that the shift checklist is done once each shift.
- Production charge taken in the 24 hours ending Friday Owl: 3.5 C
- Total production charge taken in all, end of Friday Owl: 36.1 C
Shift Summaries and BTA Reports